# A Comprehensive Framework for Reliable Legal AI: Combining Specialized Expert Systems and Adaptive Refinement

### v1

@@ 1 Introduction
In the following sections, the Literature Review (Section 2) of this paper highlights the developments in LLMs like GPT-4 for legal tasks, addressing challenges such as domain specificity and hallucinations. Section 3 details the Methodology, which integrates cutting-edge AI techniques like RAG, KGs, a MoE framework, and RLHF to deliver accurate and reliable legal analysis. This section also includes subsections for Dataset Selection, which outline the diverse datasets employed to evaluate the system. The System Architecture subsection provides a detailed description of the working of the system which employs a multi-stage workflow enhanced by KGs.
The RLHF Component (Section 8) improves system alignment with user feedback. The Results and Discussion (Section 9) showcase significant performance gains in metrics like Rouge-L and BLEU, demonstrating the framework’s strengths. Finally, the Conclusion (Section 10) highlights the system’s effectiveness and potential future enhancements, including expert module expansion and real-time updates.
We will explore the design and development of this system in detail, outlining the key challenges it addresses and the potential it offers for the future integration of AI within the legal sector.

### v2

@@ 1 Introduction
In the following sections, the paper discusses key aspects of its proposed framework. The Introduction sets the stage by highlighting challenges in integrating AI into the legal profession, such as inaccuracies and the importance of reliable systems. The Literature Review, refer Section 2, explores current AI techniques in the legal domain, focusing on large language models, domain-specific methods, and approaches to minimize hallucinations. The Methodology section, refer Section 3, presents the novel framework, detailing components like Retrieval-Augmented Generation, Knowledge Graphs, a Mixture of Experts approach, and Reinforcement Learning from Human Feedback. In Section 4 we have discussed Results and highlighted the framework’s performance improvements over existing models in legal tasks. Finally Section 5, the Conclusion, summarizes findings and suggests directions for future research in advancing reliable legal AI systems.